OpenStudy (anonymous):

Find a cubic function with the given zeros. -6, 7, -4

OpenStudy (jdoe0001):

x = -6 => x+6 = 0 => (x+6) = 0 x = 7 => x-7 = 0 => (x-7) = 0 x = -4 => x+4 = 0 => (x+4) = 0 those are your roots, multiply them to get the function/polynomial

OpenStudy (jdoe0001):

3 roots, 3rd degree polynomial, or cube function